AP Stat Unit 2 Quiz

0.0(0)
Studied by 0 people
call kaiCall Kai
learnLearn
examPractice Test
spaced repetitionSpaced Repetition
heart puzzleMatch
flashcardsFlashcards
GameKnowt Play
Card Sorting

1/38

encourage image

There's no tags or description

Looks like no tags are added yet.

39 Terms

1
New cards

short run probaiblity is

unpredictable

2
New cards

long run probability is

regular and predictable

3
New cards

Law of Large Numbers

simulated probabilities tend to get closer to the true probability as the number of trials increases

4
New cards

Guideline for interpreting probability

After Blank (how every many spins/ flips etc) the percent of times CONEXT approaches %%% in the long run

5
New cards

simulation

a way to model random events such that simulated outcomes closely watch the read world outcomes

6
New cards

statistically significant

if the number is less then 5%- means that unlikely for this event to happen purely by chance

7
New cards

sample space

list of all possible outcomes

8
New cards

P(E)=

# of outcomes/ total # of outcomes in sample space


9
New cards

compliment rule

P(AC) =1-P(A)

10
New cards

if you have P(A∩B) what do you do

multiply P(A) x P(B)

11
New cards

If you have P(A U B) what do you do

add P(A) x P(B)

12
New cards

mutually exclusive

two events that can never occur together at the same time

13
New cards

Addition

For P( A or B)=P(A) + P(B) - P(A and B)

14
New cards

The probability of any outcome or event from a random process is a number between

0 and 1 that describes the proportion of times the outcome or event would occur in a very long series of trials

15
New cards

how to perform a simulation to estimate the probability of an outcome or event

Describe how to set up and use a random process to perform one trial of the simulation. Identify what you will record at the end of each trial.

Perform many trials.

Use the results of your simulation to answer the question about the outcome or event of interest.

16
New cards

outcome

result of one trial

17
New cards

event

collection of outcomes

18
New cards

how to design a simulation

  1. assign

  2. describe one trial

  3. repeat many times

  4. conclude


19
New cards

probability of the sample space is equal to

1

20
New cards

a zero probability is

impossible

21
New cards

compliment

the probability of NOT E = 1- probability of E

22
New cards

Mutually Exclusive

can not happen at the same time (justify that a and b can both not happen at the same time)

23
New cards

In a venn diagram if there is no mutually exclusive events, do the ciricles overlap

no

24
New cards

conditional probability

p(A given B already happening)= Probability (a and b occuring at the same time/ Probability of B)

25
New cards

independent events

knowing one happened doesn’t change the probability of the other

26
New cards

two event are independent if

P(A|B) = P(A)

27
New cards

union meaning

A OR B HAPPENIGN

28
New cards

All probabilities are between

0-1 or0%-100%

29
New cards

interpreting P(A)= 0.25

if this process were repeated a larger number of times, about 25% of the repetitions would result in A

30
New cards

Estimate the P(Event)

number of times the event occured/ number of trials

31
New cards

Random outcomes have no…

memories- past results dont change the probability of the next outcome

32
New cards

If a question asks you to use results to estimate probability do

the number of times that thing occured/ the number of trials completed

33
New cards

P(at least 1)

1-P(none)

34
New cards

if A and b are not mutually exclusive, then use the

general addition rule

35
New cards

How to tell if an event is mutually exclusive

find P(A and B) if it is = 0, then it is mutually exclusive if not, then it is not mutually exclusive

36
New cards

when running a single trial, you want to get the result that

simulates the same number of trials completed in the actual trial and with the same percent as the probability

37
New cards

Whenever a question asks you to find P(A orB) what must you look for first

whether or not the events are mutually exclusive!!! if they are not, then you must use the general addition rule

38
New cards

if a question says what is the probability of both candies being one color or another color what should you do

P(A and B) P(A) X P(B)

39
New cards

to find the probability of an event happening with a given number of trials multiply

the number of trials x the predicted probability